All systems are working normally
October 23rd – 15:00 The email service fault has now been corrected and appears to be working normally. A fault had developed with the “dovecot service” (which is the program controlling incoming/outgoing emails) which required all active pop and imap connections to be “flushed” in order to restore normal service. We will continue monitoring the service.
October 22nd – 15:00 A number of users with @paston.co.uk email addresses are reporting problems receiving emails. They are able to send and able to view their email using http://webmail.paston.co.uk but the emails are not arriving on their computer. Our email provider is investigating the problem, and will update this ticket.
October 10th – 09:10 – Email services have resumed.
October 10th – 09:00 Our mail server provider is reporting a problem for customers with email addresses ending @paston.co.uk which they are investigating.
January 15th – 12:30 Our mail server provider says “We are rolling out updates and patches to all our servers. As these vulnerabilities are rated critical we will be rebooting the servers during working hours. We apologise for any issues caused by this, but as you can imagine, we need to ensure that the servers are secured from the Meltdown and Spectre vulnerabilities.”
August 10th – 17:10 – Email services have resumed.
August 10th – 15:00 – Some customers may be experiencing email problems, we are addressing this. We will update with either an ETA or resolution asap.
February 28th – 09:00 – All customers with email addresses ending @paston.co.uk will be moving to a new mail server starting at 09:00.
Because of propagation delays some customers may not get transferred immediately, but later during the day.
For many it will not be necessary to make any changes to their settings.
However customers using either post.paston.co.uk, mail.paston.co.uk or smtp.paston.co.uk as their outgoing mail server and port 2481 (in Advanced settings) need to change to cloud37.unlimitedwebhosting.co.uk as their outgoing mail server using port 25025 (in Advanced settings).
January 11th – 14:15 – 85145 spam emails were removed from the mail queue – all being sent by one particular user. The main issue here is compromised email accounts probably due to weak password sending out spam. This has a snowball effect overwhelming our server. All users with passwords that are considered “weak” by the server will be notified shortly with instructions on how to change their passwords which they should action.
January 11th – 08:00 – Rackspace are investigating an email problem that is preventing some customers from collecting email.
December 13th 11:30 – The email server is now running normally. However, Rackspace had to change some of the very weak email passwords to stop further spam arriving. If you now find your password is not being accepted please telephone Paston Chase and they will assist.
December 13th 08:30 – Some customers may still be experiencing delays receiving email. This is caused by large quantities of spam being sent out from email accounts with weak passwords. For information about how to change your password to a secure one, please click on the following link http://www.paston.co.uk/changing-email-password/
December 12th 16:00 – With nearly 100,000 “undeliverable” emails in the deferred queue on the mail server it will take sometime to fully restore normal service, but the process is continuing and the ticket will be updated when complete.
December 12th 11:15 – The problem has been isolated to one particular email account that had a very “weak” password, which had been used by spammers to send out large amounts of spam. This has now been stopped, but it will take sometime to restore normal service. All users should ensure their passwords are secure. A password should be at least 8 characters, include lower and uppercase letters and at least 1 number. Ideally you should also use special characters such as !$@&?*
December 12th 08:21 – One of our mail servers is currently being re-booted by Rackspace as it had become un-responsive, and this will effect some customers.
November 30th 12:35 – Some customers are unable to receive emails. Rackspace (the mailserver management company) are investigating. Resolved 13:42
November 23rd 07:30 – Some customers are unable to send out email. Rackspace (the mailserver management company) are investigating. Resolved 11:30
October 28th 10:10 – Rackspace (the mail server management company) notified us that there was an incoming email problem from approximately 07:50 until 10:00 which has now been cleared. However there will be some delay in email being delivered because of the backlog that built up during this period.
September 1st 2016 09:00 – We believe that the email service is now functioning normally, but Rackspace will continue to monitor the situation.
August 31st 2016 11:00 – Some customers are reporting problems with downloading large quantities of duplicated emails going back over several months. We have raised this as a problem with Rackspace (the mailserver management company) and it is being investigated.
August 26th 2016 08:30 – We think the service is now operational. All email addresses should now be functioning, although not yet fully restored with historical emails, which should complete over the weekend. Some customers may have problems sending email and these are being dealt with on a case by case basis.
August 24th 2016 10:30 – Although there are still some individual issues to be resolved, we think the service is now operational, except for the outgoing service on smtp.paston.co.uk, which is still being looked at.
August 24th 2016 08:30 – Email services are now available for @paston.co.uk email addresses.
Email that was sent to you during the downtime will have been returned to the sender as “undeliverable”.
Customers will not be able to send out through smtp.paston.co.uk – this problem is now being looked at.
August 24th 2016 01:30 – Email has been tested for domain names, excluding paston.co.uk and appear to be working. Historical emails still need to be restored for email users @paston.co.uk domain.
August 23rd 2016 14:30 – Here is an update on the restore currently in progress:
We have restored 106Gb out of 135Gb, which represents 932,119 files out of a total of 1,076,356.
We will update this status page when more information is available.
August 23rd 2016 10:30 – Work continues on restoring all the mailboxes, and we will update this status page when there is more information.
August 23rd 2016 09:30 – As the restore progresses some users may get prompted for their email username and/or password. These will not work until the system is fully restored, so please ignore.
August 23rd 2016 05:30 – Plesk control panel has been installed. The process of getting domains, users, databases, etc. back is underway. The final step is moving the existing mailbox files.
August 22nd 2016 18:20 – Work continues to keep moving the restore of the mail server forward. We are currently in this process and will update you once we have more information.
August 22nd 2016 15:30 – Hardware tests have now been completed and it has been passed back for ADC testing. We are currently in this process and will update you once we have more information.
August 22nd 2016 12:00 – ADC tests are still failing and it will be necessary to re-install the system and re-do the tests. We are currently in this process and will update you once we have more information.
August 22nd 2016 10:31 – ADC finished, however it returned few things which we need to address manually. We are currently in the process and will update you once we have more information.
August 22nd 2016 09:30 – The system checker program, ADC (Automatic Device Configuration), did not finish successfully as server could not boot properly. This has been corrected and ADC has been triggered again. This ticket will be updated as more information becomes available, at approximately 10:30am.
August 22nd 2016 07:45 – Having replaced the hardware, Rackspace are now running system checks before re-storing the applications and data. If the system checks are successful we would hope the service would be resumed at lunchtime. This ticket will be updated as more information becomes available.
August 21st 2016 21:25 – Rackspace are continuing with the restore of data from backup. This ticket with further information will be updated as more information becomes available.
August 21st 2016 14:50 – Rackspace (the company that manages our mail server) have finished the re-imaging the mailserver disks and will attempt the restore of data from backup next (3 -4 hours delay for this complete). This ticket with further information will be updated as more information becomes available.
August 21st 2016 12:40 – Rackspace (the company that manages our mail server) are continuing with re-imaging the mailserver disks and will update this ticket with further information as it becomes available.
August 21st 2016 12:10 – Rackspace (the company that manages our mail server) are re-imaging the mailserver disks, and will update this ticket at approximately 12:30pm
August 21st 2016 09:00 – The mailserver has had a file system corruption, and engineers are having to restore from a previous backup.
August 20th 2016 18:30 – Currently investigating a potential email password problem
July 21st 14:10 – Services restored and are now being monitored for any anomolies.
July 21st 13:00 – Full service restoration now expected at around 14:15.
July 21st 11:00 – Restoration of power to data center and full service expected early this afternoon.
July 21st 09:45 – Power outage at a data centre. Re-routed traffic so that internet connectivity is restored, however some customers might need to reboot their routers to restore service. Internet speeds may be slower than normal due to network congestion.
July 21st 08:14 – Broadband Engineers have identified a major non-geographic network outage affecting all connections.
May 4th 2016 08:00 – Some users maybe experiencing problems sending out email. The engineers at Rackspace (the company that handles our outgoing email) are investigating the problem, and hopefully it will be resolved shortly. Resolved 10:30.
April 7th 2016 12:45 – Currently investigating a potential email password problem, where passwords are not being recognised.
Nov 23rd 2015 17:30 – Engineers are currently investigating an issue affecting some customers using the internet. Initial investigations show a possible Distributed Denial of Service (DDoS) attack is the cause of the incident.
Resolved 20:06 – Engineers have resolved the earlier issue that was affecting internet connectivity. We are currently monitoring the network to ensure that the issue has been fully resolved.
Sept 2nd 2015 15:00 – Outgoing email services and currently not working for some customers. This problem is being investigated by our system administration services, and hopefully the problem will be resolved shortly.
Resolved 3rd Sept 10:15
May 13th 2015 16:20 – Outgoing email services and currently not working for some customers. This problem is being investigated by our system administration services, and hopefully the problem will be resolved shortly. Resolved 19:03
April 7th 2015 12:45 – Incoming and outgoing email services are currently not working for some customers. This problem is being investigated by our system administration services, and hopefully the problem will be resolved shortly. Resolved 13:10. The problem was caused by an email account sending out large amounts of spam, and blocking legitimate emails. As a precaution customers may want to resend any emails they’ve sent since 11:30
March 27th 2015 13:30 – Incoming mail services are currently not working for some customers. This problem is being investigated by our system administration services, and hopefully the problem will be resolved shortly. Resolved 15:30